OpenStudy (anonymous):

Solve for r. 11r + 10 = -7r - 19 - 11r

pooja195 (pooja195):

Add 18r to both sides. 11r+10+18r=−18r−19+18r 29r+10=−19 Subtract 10 from both sides what ru left with?

OpenStudy (mollynicole2015):

Simplifying 11r + 10 = -7r + -19 + -11r Reorder the terms: 10 + 11r = -7r + -19 + -11r Reorder the terms: 10 + 11r = -19 + -7r + -11r Combine like terms: -7r + -11r = -18r 10 + 11r = -19 + -18r Solving 10 + 11r = -19 + -18r Solving for variable 'r'. Move all terms containing r to the left, all other terms to the right. Add '18r' to each side of the equation. 10 + 11r + 18r = -19 + -18r + 18r Combine like terms: 11r + 18r = 29r 10 + 29r = -19 + -18r + 18r Combine like terms: -18r + 18r = 0 10 + 29r = -19 + 0 10 + 29r = -19 Add '-10' to each side of the equation. 10 + -10 + 29r = -19 + -10 Combine like terms: 10 + -10 = 0 0 + 29r = -19 + -10 29r = -19 + -10 Combine like terms: -19 + -10 = -29 29r = -29 Divide each side by '29'. r = -1 Simplifying r = -1 So now how do we know if it's right?

OpenStudy (anonymous):

Plug it back into the equation. 11 x -1 + 10 = -7 x -1 - 19 - 11 x -1 11 x -1 + 10 = -1 -7 x -1 - 19 - 11 x -1 = -1 -1 = -1

OpenStudy (mollynicole2015):

Good, and that means, what?

OpenStudy (anonymous):

The equation is true. :)

pooja195 (pooja195):

Sigh.

OpenStudy (mollynicole2015):

Fantastic! <3

OpenStudy (mollynicole2015):

Oh, Pooja, I didn't see you there.

OpenStudy (kalynne):

do i get a medal for this but your answer would be 65 on both cause they both equal the same thing like for example 4=4

OpenStudy (kalynne):

whoops wrong answer its 3>